OpenStudy (shamil98):

How do I find the focus and directrix of the following parabola? x = (-1/8 )y^2

OpenStudy (nincompoop):

what method you want to use?

OpenStudy (anonymous):

seperate y^2

OpenStudy (anonymous):

compare with y^2=4ax a=your focus

OpenStudy (nincompoop):

always

OpenStudy (anonymous):

medal?

OpenStudy (nincompoop):

naw... the medal goes to the person that answered first

OpenStudy (shamil98):

so y^2 = 4(-8) x? ..

OpenStudy (anonymous):

:")

OpenStudy (anonymous):

yep @shamil98

OpenStudy (shamil98):

Okay, thank you. What about the directrix?

OpenStudy (nincompoop):

x+y^2/8 = 0

OpenStudy (anonymous):

I'll give @nincompoop the oppurtunity to answer that!

OpenStudy (nincompoop):

directx is 2

OpenStudy (shamil98):

explain pls bruh

OpenStudy (shamil98):

Oh and also is the focus and directrix written in coordinates?

OpenStudy (nincompoop):

we know that parabola is y = a(x-h)^2 + k correct?

OpenStudy (shamil98):

yeah

OpenStudy (nincompoop):

then the relationships can be shown y – k = a(x – h)^2 (1/a)(y – k) = (x – h)^2 4p(y – k) = (x – h)^2 <-- conics form of the equation where 4p = 1/a

OpenStudy (shamil98):

how did you get 2 as the directrix, nin?

OpenStudy (anonymous):

oh this is still going on :O

ganeshie8 (ganeshie8):

once u have focus, and vertex, directrix goes thru the image of focus

OpenStudy (nincompoop):

p is your distance between the vertex and focus, which is also the distance between the vertex and directrix

OpenStudy (nincompoop):

my connection goes in and out

ganeshie8 (ganeshie8):